Given the system of equations.
4x+y=8
6x-9y=12
Which of the options represents the resulting equation after an equivalent expression for y is substituted into the second equation?
6x - 9(4x - 8) = 12
4x + 4x - 8 = 8
6(4x - 8) - 9y = 12
6x - 9(-4x + 8) = 12
I also do not understand how to find this, so any help would be nice on that subject

Answer by jim_thompson5910(35256)

Start with the first equation 4x+y=8.

Isolate y. In other words, get y all by itself on its own side. We do this by subtracting 4x from both sides

4x+y=8

4x+y-4x=8-4x

y=8-4x

y=-4x+8

------------------------------------------------------

The equation y=-4x+8 basically says that y is the same as -4x+8. Because of this, we can replace every 'y' in the second equation with '-4x+8'

6x-9y=12

6x-9( y )=12

6x-9( y )=12

6x-9( -4x+8 )=12 ... note how the terms in blue indicate how the replacement is happening

So the answer is choice D